About 1% of the population has a particular genetic mutation. 800 people are randomly selected. Find the standard deviation for the number of people with the genetic mutation in such groups of 800. Round your answer to three decimal places.
Answer #1

To find the standard deviation we use the formula:

s = Sqrt(n*(p)*(1-p)), where n is the sample size and p is the probability.

S = Sqrt(800*(0.01)*(1-0.01))

= Sqrt(800*(0.01)*(0.99))

=2.814